World-Wide Enterprises (WWE) is planning to enter into the business of PC manufacturing. IBM is a typical firm in this industry. and can be used as a reference. WWE has a DE of 0.7, IBM has a DE of 0.4. Borrowing rate for WWE is 12 %, while borrowing rate for IBM is 10%. Given Market risk premium is 8 %, risk-free rate is 5%, and corporate tax rate 34%, what is the appropriate discount rate for WWE to use for its PC venture?
Answer #1

Using publicly available data, we know beta of IBM is 1.87
Now this is levered beta and we need to find unlevered beta

Bu = Bl/(1+(1-t)*D/E)

Bu = 1.87/(1+(1-0.34)*0.4) = 1.479

For WWE, we can now find levered beta using the D/E of WWE

Bl = Bu*(1+(1-t)*D/E)
Bl = 1.479*(1+(1-0.34)*0.7) = 2.1623

So Ke = 5% + 2.1623*8% = 22.30%

Kd = 12%

Now D/E = 0.7 So E/(D+E) = 1/1.7 = 0.588

D/D+E = 0.412

WACC = 0.412*12%*(1-0.34) + 0.588 * 22.30% = 16.38%